OpenStudy (anonymous):

f(x)=5 x -4 and g(x)=3 - x^2 find f(g(0))

OpenStudy (lanhikari22):

well it's up to you, whether you solve g(0) first and replace the answer with x in f(x) or whether you put them all together as (f o g) and find (f o g)(0) first case g(0) = 3 - (0)^2 -> g(0) = 3, ok put that in f(x) f(g(x)) = f(3) = 5(3) -4 = 15 - 4 = 11 that or just do (f o g)(0) = 5 ( 3 - x^2) -4 = 15 - 5x^2 - 4 since x= 0 theeen 15 - 0 - 4 = 15 - 4 = 11 aaand that's it. hope i helped.

OpenStudy (anonymous):

how would i fing g(f(0))

OpenStudy (aum):

To find g(f(0)) first evaluate the innermost function. That is, find f(0). You will get a number Put that number in g( number ) and find g( number ).